# Efficiently learning depth-3 circuits via quantum agnostic boosting

### v2

@@ 2.4.3 Quantum agnostic learning
The sample complexity of learning is the total number of copies used by the algorithm to satisfy the above guarantee and the time complexity is the total time. We say an algorithm is sample and time efficient these complexities scale polynomial in $n,1/\varepsilon$ and the the description size of the class. If the learner outputs $|\psi^{\prime}\rangle\in\mathcal{S}$, then the learner is called proper, else its an improper learner. As far as we are aware, there are only a handful of works that have considered quantum state agnostic learning [CGYZ25, BO21, BBK+25] wherein they considered interesting classes of states such as stabilizer states, product states, matrix product states and proved results for this model. As we mentioned in the introduction, in this work, we will be concerned with the concept class of states being
@@ 3.1 Useful subroutines and lemmas
For a set of $k$ states $\{|\phi_{i}\rangle\}_{i\in[k]}$, let its span be denoted as $T=\mathsf{span}(\{|\phi_{i}\rangle\}_{i\in[k]})$. Let T be a projection onto $T$. The projection of $|\psi\rangle$ onto $T$ is given by
@@
Lemma 3.6 (Clifford synthesis [DDM03, PMH03]).
@@ 3.2 Algorithm
$\langle\psi|\mathbb{I}-{}_{T(t)}|\psi\rangle=\langle\psi|(\mathbb{I}-\sumop\displaylimits_{i\in[t]}|\phi_{i}\rangle\langle\phi_{i}|)|\psi\rangle=1-\sumop\displaylimits_{i\in[t]}|\langle\psi|\phi_{i}\rangle|^{2}=1-\sumop\displaylimits_{i\in[t]}\beta_{i}^{2}=|\alpha_{t+1}|^{2}\geq\varepsilon,$
@@
In the previous stage we learned the new parity function, but in order to update our current state $|\widehat{\phi}^{(t)}\rangle=\sumop\displaylimits_{i\in[t]}\beta_{i}|\phi_{i}\rangle$ to $|\widehat{\phi}^{(t+1)}\rangle=\sumop\displaylimits_{i\in[t+1]}\beta_{i}|\phi_{i}\rangle$, we also need to learn the coefficient $\beta_{t+1}$. To do so, one could compute $\beta_{t+1}=\langle\phi_{t+1}|\psi\rangle$ via the Hadamard test using the state preparation unitaries (and their controlled versions) of $|\phi_{t+1}\rangle$ and $|\psi\rangle$. However, we show that we can avoid the need for state preparation unitary access and accomplish the task of agnostic learning by determining $_{T(t)}|\psi\rangle$ up to a global phase, using only copies of $|\psi\rangle$. This is formally stated in Lemma 3.11. We are now ready to present the algorithm.
@@ 3.3 Structure learning
$1\geq\mbox{$\parallel{|\psi_{1}\rangle}\parallel$}_{2}^{2}-\mbox{$\parallel{|\psi_{\kappa}\rangle}\parallel$}_{2}^{2}=\sumop\displaylimits_{t=1}^{\kappa}\mbox{$\parallel{|\psi_{t}\rangle}\parallel$}_{2}^{2}-\mbox{$\parallel{|\psi_{t+1}\rangle}\parallel$}_{2}^{2}\geq\kappa\varepsilon_{s}\eta(\varepsilon)/4\implies\kappa\leq 4/(\varepsilon_{s}\eta(\varepsilon)),$
@@ 3.4 Parameter learning
$\left|\langle\psi\Big|\Big(\sumop\displaylimits_{i=1}^{k}\widehat{\beta}_{i}|\phi_{i}\rangle\Big)\right|^{2}\geq|\langle\psi|({}_{T}|\psi\rangle)\rangle|^{2}-\gamma,$ (20)
@@
$|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}c_{i}|\phi_{i}\rangle,$ (22)
@@
$|\langle\psi|\widehat{\phi}\rangle|^{2}=\frac{\left|\langle\psi\Big|\Big(\sumop\displaylimits_{i=1}^{k}\widehat{\beta}_{i}|\phi_{i}\rangle\Big)\right|^{2}}{\mbox{$\parallel{\widehat{\beta}}\parallel$}_{2}^{2}}$ $\geq\frac{|\langle\psi|({}_{T}|\psi\rangle)\rangle|^{2}-\gamma}{|\langle\psi|({}_{T}|\psi\rangle)\rangle|^{2}+\gamma}$
@@
Let us define the states $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{k}\widehat{\beta}_{i}|\chi_{i}\rangle$ and $|\widetilde{\chi}\rangle=\sumop\displaylimits_{i=1}^{k}\widetilde{\beta}_{i}|\phi_{i}\rangle$, which are not necessarily normalized. Eq. (56) then implies
$|\langle\psi|\widehat{\phi}\rangle-\langle\psi|\widetilde{\phi}\rangle|=\Big|\sumop\displaylimits_{i=1}^{k}(\widehat{\beta}_{i}-\widetilde{\beta}_{i})\langle\psi|\phi_{i}\rangle\Big|\leq\sumop\displaylimits_{i=1}^{k}|\widehat{\beta}_{i}-\widetilde{\beta}_{i}|\leq\varepsilon/3.$ (57)
@@
where we have used in the second line that $|\widetilde{\beta}_{i}|=|\beta_{i}|=|\langle\psi|\phi_{i}\rangle|\leq 1$ and $\sumop\displaylimits_{i}|a_{i}|^{2}\leq(\sumop\displaylimits_{i}|a_{i}|)^{2}$. In the third line, we used Eq. (56) and noted that $\sumop\displaylimits_{i=1}^{k}|\widetilde{\beta_{i}}|^{2}=\sumop\displaylimits_{i=1}^{k}|\beta_{i}|^{2}=|\langle\psi|({}_{T}|\psi\rangle)|^{2}$.
@@ 4.2 Agnostic learning decision trees
with probability $\geq 1-\delta$. Moreover, $|\widehat{\phi}\rangle$ can be expressed as $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\phi_{i}\rangle$ with $\beta\in{\mathcal{B}}_{\infty}^{k}$ being coefficients corresponding to $|\phi_{i}\rangle$, which are parities, and $\kappa=\operatorname{poly}(s/\varepsilon)$.
@@
where $\kappa=\operatorname{poly}(2^{k}/\varepsilon)$ and $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\phi_{i}\rangle$ with $\beta\in{\mathcal{B}}_{\infty}^{k}$ being coefficients for the parities $|\phi_{i}\rangle$.
@@ 4.3 Agnostic learning DNFs
where $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\phi_{i}\rangle$ with coefficients $\beta\in{\mathcal{B}}_{\infty}^{k}$ for the parities $|\phi_{i}\rangle$ and $\kappa=\operatorname{poly}((s/\varepsilon)^{\log\log(s/\varepsilon)\cdot\log(1/\varepsilon)})$.
@@ 4.4.2 Learning algorithm
$\operatorname{poly}(n,m,(s/\varepsilon)^{(\log\log s/\varepsilon)\cdot\log 1/\varepsilon},\log(1/\delta)).$
To prove this, we will show that applying improper agnostic learning of ${\cal S}_{{\mathcal{C}}_{\textsf{DNF}(s)}}$ when the input state $|\psi_{f}\rangle$ is promised to a phase state corresponding to a depth-$3$ circuit, accomplishes the task of state tomography. To that end, we will revisit the quantum agnostic boosting argument in Section 3 when the goal is to do well against the class ${\mathcal{C}}_{\textsf{DNF}(s)}$. For the argument below, we now specialize our discussion to when the input is a phase state $|\psi_{f}\rangle$ since we are in the quantum PAC setting. Recall that in the agnostic boosting algorithm (Algorithm 1), we had a running estimate $|\widehat{\phi}^{(t)}\rangle$ defined as the projection of the input state $|\psi_{f}\rangle$ on to the set of basis states $\{|\phi_{i}\rangle\}_{i\in[t]}$ learned so far:
@@
Suppose as part of structure learning of the agnostic boosting algorithm (Algorithm 1), we have carried out t1t-1 iterations so far and learned the parity states $\{|\chi_{i}\rangle\}_{i\in[t-1]}$. Then, the residual quantum state $|\psi_{t}\rangle$ in the $t$th iteration can be expressed as
$|\psi_{t}\rangle=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-{}_{T(t-1)}|\psi_{f}\rangle\Big)=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}|\chi_{i}\rangle\Big),$
where $T(t-1)=\mathrm{span}(\{|\phi_{i}\rangle\}_{i\in[t-1]})$, $\beta_{i}=\langle\chi_{i}|\psi_{f}\rangle$ (Eq. (8)) and $\alpha_{t}=\sqrt{1-\sumop\displaylimits_{i=1}^{t-1}|\beta_{i}|^{2}}$ (Fact 3.3). Let us denote the states
@@
$D_{f,t}(x)=\Big|f(x)-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x)\Big|\cdot 2^{-n}\cdot{}^{-1}$,
@@
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left|f(x)-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x)\right|\cdot f(x)\cdot c(x)$
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left(f(x)-\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x)\right)\cdot c(x)$
@@
where the third equality used the fact that, for $a\in\{-1,1\}$ and $b\in[-1,1]$, we have that $|a-b|=a\cdot(a-b)$. [^16] To see this note that if $a=1$ then $a\geq b$ and $|a-b|=a-b=1-b=a\cdot(a-b).$ If a=1,a=-1, then $a\leq b$ and $|a-b|=b-a=b+1=a\cdot(a-b).$ One can then use Lemma 4.12 to show that
@@
Let $f$ be as in the theorem statement. We will employ the quantum agnostic boosting algorithm (Algorithm 1) on input of copies of $|\psi_{f}\rangle$ against the class ${\cal C}_{\textsf{DNF}(s)}$ and use the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9.171717Our notation here is similar to the one in Section 3. Suppose we are at the end of the $t$th iteration of the agnostic boosting algorithm. We first observe that by Claim 4.15, if $\alpha_{t+1}^{2}\geq\varepsilon$, then ${\cal F}_{{\mathcal{C}}_{\textsf{DNF}(s)}}(|\psi_{t+1}\rangle)\geq\varepsilon/(4m^{2})$. If we were to then use ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9, we are guaranteed to learn a parity state $|\phi_{t+1}\rangle$ that satisfies $|\langle\chi_{t+1}|\psi_{t+1}\rangle|^{2}\geq\varepsilon/(4m^{2}s^{\star})$ where we have denoted $s^{\star}=(s/\varepsilon)^{(\log\log s/\varepsilon)\cdot\log 1/\varepsilon}$. Note that the function $\varepsilon/s^{*}$ is an increasing function of $\varepsilon$ (as required by the boosting algorithm). Thus, we no longer need to check the fidelity of the residual state with ${\cal C}_{\textsf{DNF}(s)}$. We now present the simplified stage $1$ of the algorithm in Algorithm 3
@@
where the second inequality used that $\mathop{\mathbb{E}}_{x}[h(x)^{2}]=\sumop\displaylimits_{i}\beta_{i}^{2}\leq 1$ and the assumption of this case that $\mathop{\mathbb{E}}_{x}[f(x)h(x)]\geq\sqrt{1-\varepsilon}\geq 1-\varepsilon/2$. Hence the Boolean function $g$ is an $\varepsilon$-approximator for the unknown $f\in\textsf{TAC}^{0}_{2}$ and we are done. Finally note that the learning algorithm knows explicitly $|\widehat{\phi}\rangle$, so it can output $g$ as well. The main contribution to time complexity is due to running Algorithm 3 and utilizing Theorem 3.12. The overall time complexity is $\operatorname{poly}(n,t,s^{\star})$. ∎

### v3

@@ 2.4.3 Quantum agnostic learning
The sample complexity of learning is the total number of copies used by the algorithm to satisfy the above guarantee and the time complexity is the total time. We say an algorithm is sample and time efficient these complexities scale polynomial in $n,1/\varepsilon$ and the the description size of the class. If the learner outputs $|\psi^{\prime}\rangle\in\mathcal{S}$, then the learner is called proper, else its an improper learner. As far as we are aware, there are only a handful of works that have considered quantum state agnostic learning [CGY+25, BO21, BBK+25] wherein they considered interesting classes of states such as stabilizer states, product states, matrix product states and proved results for this model. As we mentioned in the introduction, in this work, we will be concerned with the concept class of states being
@@ 3.1 Useful subroutines and lemmas
For a set of $k$ states $\{|\chi_{i}\rangle\}_{i\in[k]}$, let its span be denoted as $T=\mathsf{span}(\{|\chi_{i}\rangle\}_{i\in[k]})$. Let T be a projection onto $T$. The projection of $|\psi\rangle$ onto $T$ is given by
@@
Lemma 3.6 (Clifford synthesis [DD03, PMH03]).
@@ 3.2 Algorithm
$\langle\psi|\mathbb{I}-{}_{T(t)}|\psi\rangle=\langle\psi|(\mathbb{I}-\sumop\displaylimits_{i\in[t]}|\chi_{i}\rangle\langle\chi_{i}|)|\psi\rangle=1-\sumop\displaylimits_{i\in[t]}|\langle\psi|\chi_{i}\rangle|^{2}=1-\sumop\displaylimits_{i\in[t]}\beta_{i}^{2}=|\alpha_{t+1}|^{2}\geq\varepsilon,$
@@
In the previous stage we learned the new parity function, but in order to update our current state $|\widehat{\phi}^{(t)}\rangle=\sumop\displaylimits_{i\in[t]}\beta_{i}|\chi_{i}\rangle$ to $|\widehat{\phi}^{(t+1)}\rangle=\sumop\displaylimits_{i\in[t+1]}\beta_{i}|\chi_{i}\rangle$, we also need to learn the coefficient $\beta_{t+1}$. To do so, one could compute $\beta_{t+1}=\langle\phi_{t+1}|\psi\rangle$ via the Hadamard test using the state preparation unitaries (and their controlled versions) of $|\phi_{t+1}\rangle$ and $|\psi\rangle$. However, we show that we can avoid the need for state preparation unitary access and accomplish the task of agnostic learning by determining $_{T(t)}|\psi\rangle$ up to a global phase, using only copies of $|\psi\rangle$. This is formally stated in Lemma 3.11. We are now ready to present the algorithm.
@@ 3.3 Structure learning
$1\geq\mbox{$\parallel{|\psi_{1}\rangle}\parallel$}_{2}^{2}-\mbox{$\parallel{|\psi_{\kappa+1}\rangle}\parallel$}_{2}^{2}=\sumop\displaylimits_{t=1}^{\kappa}\mbox{$\parallel{|\psi_{t}\rangle}\parallel$}_{2}^{2}-\mbox{$\parallel{|\psi_{t+1}\rangle}\parallel$}_{2}^{2}\geq\kappa\varepsilon_{s}\eta(\varepsilon_{s})/4\implies\kappa\leq 4/(\varepsilon_{s}\eta(\varepsilon_{s})),$
@@ 3.4 Parameter learning
$\left|\langle\psi\Big|\Big(\sumop\displaylimits_{i=1}^{k}\widehat{\beta}_{i}|\chi_{i}\rangle\Big)\right|^{2}\geq|\langle\psi|({}_{T}|\psi\rangle)\rangle|^{2}-\gamma,$ (20)
@@
$|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}c_{i}|\chi_{i}\rangle,$ (22)
@@
$|\langle\psi|\widehat{\phi}\rangle|^{2}=\frac{\left|\langle\psi\Big|\Big(\sumop\displaylimits_{i=1}^{k}\widehat{\beta}_{i}|\chi_{i}\rangle\Big)\right|^{2}}{\mbox{$\parallel{\widehat{\beta}}\parallel$}_{2}^{2}}$ $\geq\frac{|\langle\psi|({}_{T}|\psi\rangle)\rangle|^{2}-\gamma}{|\langle\psi|({}_{T}|\psi\rangle)\rangle|^{2}+\gamma}$
@@
Let us define the states $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{k}\widehat{\beta}_{i}|\chi_{i}\rangle$ and $|\widetilde{\chi}\rangle=\sumop\displaylimits_{i=1}^{k}\widetilde{\beta}_{i}|\chi_{i}\rangle$, which are not necessarily normalized. Eq. (56) then implies
$|\langle\psi|\widehat{\phi}\rangle-\langle\psi|\widetilde{\phi}\rangle|=\Big|\sumop\displaylimits_{i=1}^{k}(\widehat{\beta}_{i}-\widetilde{\beta}_{i})\langle\psi|\chi_{i}\rangle\Big|\leq\sumop\displaylimits_{i=1}^{k}|\widehat{\beta}_{i}-\widetilde{\beta}_{i}|\leq\varepsilon/3.$ (57)
@@
where we have used in the second line that $|\widetilde{\beta}_{i}|=|\beta_{i}|=|\langle\psi|\chi_{i}\rangle|\leq 1$ and $\sumop\displaylimits_{i}|a_{i}|^{2}\leq(\sumop\displaylimits_{i}|a_{i}|)^{2}$. In the third line, we used Eq. (56) and noted that $\sumop\displaylimits_{i=1}^{k}|\widetilde{\beta_{i}}|^{2}=\sumop\displaylimits_{i=1}^{k}|\beta_{i}|^{2}=|\langle\psi|({}_{T}|\psi\rangle)|^{2}$.
@@ 4.2 Agnostic learning decision trees
with probability $\geq 1-\delta$. Moreover, $|\widehat{\phi}\rangle$ can be expressed as $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\chi_{i}\rangle$ with $\beta\in{\mathcal{B}}_{\infty}^{k}$ being coefficients corresponding to $|\chi_{i}\rangle$, which are parities, and $\kappa=\operatorname{poly}(s/\varepsilon)$.
@@
where $\kappa=\operatorname{poly}(2^{k}/\varepsilon)$ and $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\chi_{i}\rangle$ with $\beta\in{\mathcal{B}}_{\infty}^{k}$ being coefficients for the parities $|\chi_{i}\rangle$.
@@ 4.3 Agnostic learning DNFs
where $|\widehat{\phi}\rangle=\sumop\displaylimits_{i=1}^{\kappa}\beta_{i}|\chi_{i}\rangle$ with coefficients $\beta\in{\mathcal{B}}_{\infty}^{k}$ for the parities $|\chi_{i}\rangle$ and $\kappa=\operatorname{poly}((s/\varepsilon)^{\log\log(s/\varepsilon)\cdot\log(1/\varepsilon)})$.
@@ 4.4.2 Learning algorithm
$\operatorname{poly}(n,m,(s/\varepsilon)^{(\log s/\varepsilon\cdot\log\log s/\varepsilon)},\log(1/\delta)).$
To prove this, we will show that applying improper agnostic learning of ${\cal S}_{{\mathcal{C}}_{\textsf{DNF}(s)}}$ when the input state $|\psi_{f}\rangle$ is promised to a phase state corresponding to a depth-$3$ circuit, accomplishes the task of state tomography. To that end, we will revisit the quantum agnostic boosting argument in Section 3 when the goal is to do well against the class ${\mathcal{C}}_{\textsf{DNF}(s)}$. For the argument below, we now specialize our discussion to when the input is a phase state $|\psi_{f}\rangle$ since we are in the quantum PAC setting. Recall that in the agnostic boosting algorithm (Algorithm 1), we had a running estimate $|\widehat{\phi}^{(t)}\rangle$ defined as the projection of the input state $|\psi_{f}\rangle$ on to the set of basis states $\{|\chi_{i}\rangle\}_{i\in[t]}$ learned so far:
@@
Suppose as part of structure learning of the agnostic boosting algorithm (Algorithm 1), we have carried out t1t-1 iterations so far and learned the parity states $\{|\chi_{i}\rangle\}_{i\in[t-1]}$. Recall that until the t1t-1th iteration, the learner has obtained $\beta_{1},\ldots,\beta_{t}$ and also the functions $\{\chi_{i}\}_{i\in[t]}$. Given explicit descriptions of these, the learner can implement the map
$O_{t-1}:|x\rangle|0\rangle\mapsto|x\rangle|\textsf{ceil}\Big(\sumop\displaylimits_{i\in[t-1]}\beta_{i}\chi_{i}(x)\Big)\rangle,$
where $\textsf{ceil}(a)=a$ if $a\in[-1,1]$, $\textsf{ceil}(a)=-1$ if $a\leq-1$ and $\textsf{ceil}(a)=1$ if $a\geq 1$. The oracle can be implemented in time $O(t)$ by basic arithematic operations. Define $g(x)=\textsf{ceil}(\sumop\displaylimits_{i}\beta_{i}\chi_{i}(x))$ and $|\psi_{g}\rangle=\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}g(x)|x\rangle$ (which can be prepared by making $O(1)$ calls to the oracle $O_{t-1})$. Defining $_{g,t}=|\psi_{g}\rangle\langle\psi_{g}|$. Now measuring $|\psi_{f}\rangle$ in the basis $\{{}_{g},\mathbb{I}-{}_{g}\}$, the post measurement state if we obtain the second outcome is given by
$|\psi_{t}\rangle=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-{}_{g,t}|\psi_{f}\rangle\Big)=\frac{1}{\alpha_{t}}\Big(|\psi_{f}\rangle-\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}g(x)|x\rangle\Big),$
where $T(t-1)=\mathrm{span}(\{|\chi_{i}\rangle\}_{i\in[t-1]})$, $\beta_{i}=\langle\chi_{i}|\psi_{f}\rangle$ (Eq. (8)) and $\alpha_{t}=\sqrt{1-\sumop\displaylimits_{i=1}^{t-1}|\beta_{i}|^{2}}$ (Fact 3.3). Let us denote the states
@@
$D_{f,t}(x)=\Big|f(x)-\textsf{ceil}(\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x))\Big|\cdot 2^{-n}\cdot{}^{-1}$,
@@
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left|f(x)-\textsf{ceil}(\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x))\right|\cdot f(x)\cdot c(x)$
$=2^{-n}\cdot{}^{-1}\cdot\sumop\displaylimits_{x}\left(f(x)-\textsf{ceil}(\sumop\displaylimits_{i=1}^{t-1}\beta_{i}\chi_{i}(x))\right)\cdot c(x)$
@@
where the third equality used the fact that, for $a\in\{-1,1\}$ and $b\in[-1,1]$, we have that $|a-b|=a\cdot(a-b)$. [^16] To see this: if $a=1$ then $a\geq b$ and $|a-b|=a-b=1-b=a\cdot(a-b).$ If a=1,a=-1, then $a\leq b$ and $|a-b|=b-a=b+1=a\cdot(a-b).$ One can then use Lemma 4.12 to show that
@@
Let $f$ be as in the theorem statement. We will employ the quantum agnostic boosting algorithm (Algorithm 1) on input of copies of $|\psi_{f}\rangle$ against the class ${\cal C}_{\textsf{DNF}(s)}$ and use the weak agnostic learner ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9.171717Our notation here is similar to the one in Section 3. Suppose we are at the end of the $t$th iteration of the agnostic boosting algorithm. We first observe that by Claim 4.15, if $\alpha_{t+1}^{2}\geq\varepsilon$, then ${\cal F}_{{\mathcal{C}}_{\textsf{DNF}(s)}}(|\psi_{t+1}\rangle)\geq\varepsilon/(4m^{2})$. If we were to then use ${\mathcal{A}}_{\mathsf{WAL}}$ of Lemma 4.9, we are guaranteed to learn a parity state $|\chi_{t+1}\rangle$ that satisfies $|\langle\chi_{t+1}|\psi_{t+1}\rangle|^{2}\geq\varepsilon/(4m^{2}s^{\star})$ where we have denoted $s^{\star}=(s/\varepsilon)^{(\log\log s/\varepsilon)\cdot\log 1/\varepsilon}$. Note that the function $\varepsilon/s^{*}$ is an increasing function of $\varepsilon$ (as required by the boosting algorithm). Thus, we no longer need to check the fidelity of the residual state with ${\cal C}_{\textsf{DNF}(s)}$. We now present the simplified stage $1$ of the algorithm in Algorithm 3
@@
where the second inequality used that $\mathop{\mathbb{E}}_{x}[h(x)^{2}]=\sumop\displaylimits_{i}\beta_{i}^{2}\leq 1$ and the assumption of this case that $\mathop{\mathbb{E}}_{x}[f(x)h(x)]\geq\sqrt{1-\varepsilon}\geq 1-\varepsilon/2$. Hence the Boolean function $g$ is an $\varepsilon$-approximator for the unknown $f\in\textsf{TAC}^{0}_{2}$ and we are done. Finally note that the learning algorithm knows explicitly $|\widehat{\phi}\rangle$, so it can output $g$ as well. The main contribution to time complexity is due to running Algorithm 3 and utilizing Theorem 3.12. ∎